OpenStudy (anonymous):

Matt scored 8 more points than Ted in the basketball game, and their scores together made up a quarter of the team's total scoring, which was 112. Which equation represents this scenario? A: 4(8 + x) = 112 B: 4(2x + 8) = 112 C: 112/4 = 8 + x D: x = 112/(4 x 8)

OpenStudy (amistre64):

112/4 = T + T+8

OpenStudy (amistre64):

28 = 2T + 8 20 = 2T T = 10 and M = 18

OpenStudy (amistre64):

B...its B... I see it...it BBBB

OpenStudy (anonymous):

not looking for a sloution,,,I know the answer is B just don't know why

OpenStudy (amistre64):

We know that a total of 112 point were scored; and that M and T totaled together is 1/4 of that total: 112/4 = M + T We know that M scored 8 more points than T which gives us: 112/4 = T + T + 8 Add the Ts to get: 112/4 = 2T + 8 T is just another name for "x" so we can replace it: 112/4 = (2x + 8) We multiply both sides by "4" to get rid of the fraction thingy on the LHS: 122 = 4(2x + 8)

OpenStudy (amistre64):

and we learn how to type so that everything matches up......

OpenStudy (anonymous):

Thank you!